# Bounding Causal Effects with Leaky Instruments

### v1

@@ 5 Related Work
Violations of the exclusion restriction are well-documented in genetics [Hemani2018] and econometrics [Berkowitz2008]. One strategy for estimating causal effects in such settings is to permit a large number of potentially invalid instruments under the assumption that their average bias will tend to zero in the limit [Bowden2015, Kolesar2015]. With weak monotonicity constraints, these approaches can also provide nonparametric bounds on local average treatment effects [Flores2013].
@@
Another family of methods starts from the assumption that some proportion of candidate instruments are valid and uses statistical procedures to focus on just those variables that satisfy (A1)-(A3). This can be achieved, for instance, via goodness of fit tests [chu_semi_instrumental]; $L_{1}$-penalized regression for feature selection [kang2016, Guo2018, Windmeijer2019]; independence tests for collider bias [kang2020]; or modal validity assumptions in linear [Hartwig2017] and nonlinear [hartford2021] IV models. Alternatively, data from multiple instruments can be pooled into a single variable using dimensionality reduction techniques [kuang2020].
@@
There is a substantial literature on Bayesian approaches to causal inference in IV settings. Lenkoski2014 use Bayesian model averaging to select IVs based on the strength of their association with the treatment, as codified by the relevance criterion (A1). Shapland2019 extend this method to account for linkage disequilibrium in Mendelian randomization experiments, which violate the no confounding condition (A2). More recently, several authors have proposed spike-and-slab priors to select genetic variants in the face of horizontal pleiotropy [bucur2020, Gkatzionis2021], thereby addressing (A3).
@@
Conley2012 propose ATE bounding methods given various kinds of prior information on leaky coefficients $\bm{\gamma}$, including a range restriction and a (non-uniform) prior distribution. For the former, they return a union of confidence intervals resulting from a grid of possible values for $\bm{\gamma}$. This method scales exponentially with $d_{\bm{Z}}$ and is potentially conservative as grid resolution grows finer. By contrast, our convex optimization approach is an efficient one-shot procedure that provides provably sharp bounds—in closed form, for the $L_{2}$ case. Their Bayesian proposal is similar to the method we compare against in Fig. 6.
@@
The literature on tetrad constraints in linear SEMs goes back over a century [spearman1904], although the method was revived and refined following the publication of Spirtes2000’s tetrad representation theorem. Our exclusion test builds on generalized results developed by numerous authors [shafer_tetrads, sullivant2010, spirtes2013], although we are to our knowledge the first to propose a Monte Carlo inference procedure for testing such claims.
@@
Partial identification intervals for counterfactual quantities can be computed exactly for discrete variables by formulating the problem as a polynomial program [zhang2022, duarte2023]. Though continuous data can always in principle be discretized with arbitrary precision, this quickly becomes intractable in the response function framework, as it leads to an exponential explosion of parameters. Some continuous alternatives have been proposed with applications to IV models [kilbertus2020class, hu2021, padh_stochastic2023]. However, the neural architectures underlying these models can be notoriously unstable, and are ill-suited to the linear SEM setting, which is standard in much biological and econometric research.
@@ 6 Discussion
We have assumed in this paper that treatment effects are homogeneous throughout the population. However, a great deal of recent literature in causal machine learning has focused on heterogeneous treatment effects, where potential outcomes are presumed to vary as a function of pre-treatment covariates [Chernozhukov2018, Kunzel2019, Nie2021]. Some authors have brought this framework into IV models, showing that tighter ATE bounds are possible with the help of observed confounders [cai2007, hartford2021, kennedy2023]. Future work will consider conditional bounding methods, where extrema for $\theta$ may depend on instruments and/or other features causally antecedent to $X$.
A final note is that our method relies on user specification of the hyperparameter $\tau$. Ill-chosen thresholds may lead to issues, either in the form of overly conservative bounds (if $\tau$ is too high) or no bounds at all (if $\tau$ is below the minimum consistent with the data). In the worst case, invalid bounds will result from selecting a threshold that is high enough to satisfy the partial identification criterion but underestimates the true value of $\lVert\bm{\gamma}\rVert_{p}$. However, we emphasize that $\tau$-exclusion is a strictly weaker assumption than the classical exclusion criterion (A3), which is widely applied—rightly or wrongly—in IV analyses.
In many cases, setting $\tau=0$—i.e., assuming perfectly exclusive instrumental variables—is an essentially arbitrary choice, and a tempting one given the veneer of certainty provided by obtaining a point estimate for the ATE. If nothing else, we hope to make practitioners think twice before falling back on this familiar default. Background knowledge is regularly used to guide hyperparameter selection, and it is reasonable to assume that in many applications practitioners will have an a priori sense of how much information leakage is likely between $\bm{Z}$ and $Y$. In particular, we can interpret this as a type of sensitivity analysis that asks how large $\tau$ can be such that the bounds exclude zero or effects of particular magnitude, and inquire with a practitioner whether larger values of $\tau$ are scientifically plausible.
@@ A.1 Proof of Lemmas 1 and 2
Consider Eqs. 1, 2, and 5, which define our model. Evaluating the product of $X$ and $\bm{Z}$ gives a relationship between covariances:
@@ A.3 Proof of Lemma 4
where $\theta^{*},\gamma^{*}$ denote the true unobservable parameters. By the convexity of the norm, the solutions to Eq. 14 for any $\tau\in[\check{\tau}_{p},\tau^{*}_{p})$ will fail to capture the true ATE, as the resulting horizontal line lies below the point $(\lVert\bm{\gamma}^{*}\rVert_{p},\theta^{*})$. Even a $\bm{\Sigma}$-oracle—who, recall, has access to the population covariance matrix but not the latent parameters $\theta^{*},\bm{\gamma}^{*}$—will return invalid bounds if queried with a threshold in this half-closed interval. For this reason, we call this band the error region.
@@
We have thus far assumed that the oracle threshold strictly exceeds the theoretical minimum, but these parameters may coincide. If $\tau^{*}_{p}=\check{\tau}_{p}$, then the error region is empty and all identifiable bounds are valid. Moreover, if $g_{p}$ attains a unique minimum—as it must for all strictly convex norms, i.e. $p\in(1,\infty)$—then there exists just a single solution to Eq. 14 for $\tau=\tau^{*}_{p}=\check{\tau}_{p}$. In this case, lower and upper bounds for the ATE coincide and the causal parameter is point identified as $\theta^{*}=\check{\theta}_{p}$. Note that this fortuitous circumstance occurs with Lebesgue measure zero, as it imposes a nontrivial polynomial constraint on covariance parameters [okamoto1973].
@@
For every $P_{\bm{\Sigma}}\in\mathcal{M}$, there exists some nearest null neighbor $Q_{\bm{\Sigma}}\in\mathcal{M}_{0}$ (not necessarily unique) satisfying
@@
for all $u\in[0,1]$ [Lehmann2005]. In other words, $p_{\text{MC}}$ is uniformly distributed under $H_{0}$, as desired, and the Monte Carlo null distribution has converged on the target $G_{n}^{\bm{\Sigma}^{0}}$. We add one to the numerator and denominator as a necessary finite sample adjustment.
@@
It may not be immediately clear that bootstrapping is appropriate for ATE bounds. After all, it is well known that the bootstrap cannot provide a valid sampling distribution for fixed order statistics such as ranks, or a target parameter that lies on the boundary of the parameter space [andrews2000]. But though we refer to our bounds as “min” and “max” solutions, that does not mean they are calculated via fixed order statistics. On the contrary, each represents a continuous solution to a differentiable optimization task, not the smallest or largest element in a discrete set. The only boundary condition for our target parameters is $\theta^{-}\leq\theta^{+}$, which automatically holds under the partial identifiability criterion $\tau\geq\check{\tau}_{p}$. As our estimator is undefined for samples that violate the criterion, our sampling distribution is always conditioned on this event.
In general, any statistic that is a differentiable function of sample moments admits an Edgeworth expansion and can therefore have its distribution consistently estimated via bootstrap resampling [hall1992, davison_hinkley]. Recall that our ATE bounds represent the intersection of (a) a $\tau$-feasible region that is fixed a priori; and (b) the $L_{p}$ norm of $\bm{\gamma}=\bm{\alpha}-\theta\bm{\beta}$, where the latter two vectors are defined as dot products of covariance parameters. Resulting bounds vary smoothly under resampling, since $\bm{\alpha}$ and $\bm{\beta}$ are differentiable with respect to $\bm{\Sigma}$. Though more generalized relaxations of the exclusion criterion may introduce discontinuities or other issues, our formulation of $\tau$-exclusion poses no such difficulties. The resulting bootstrap distributions are asymptotically valid and practically useful, providing statistical inference without any parametric assumptions.
Of course, it is perfectly possible that some bootstrap samples may have to be discarded if the intersection of regions (a) and (b) is empty. In such cases, we simply restrict attention to those bootstraps that satisfy the partial identifiability criterion $\tau>\check{\tau}_{p}$, which should represent a non-negligible proportion of all bootstraps if the inequality is satisfied in the original dataset. This procedure is akin to sampling under a feasibility condition, and requires no extra steps to maintain bootstrap consistency, as in andrews2000 and ramsahai_likelihood_2011.
@@ B.3 Bayesian baseline
$\begin{array}[]{rcl}\bm{\Sigma}(\Theta)_{\bm{zz}}&:=&\ \bm{\Sigma}_{\bm{zz}}\\ \bm{\Sigma}(\Theta)_{\bm{z}x}&:=&\ \bm{\Sigma}_{\bm{zz}}\beta\\ \bm{\Sigma}(\Theta)_{xx}&:=&\bm{\beta}\cdot\bm{\Sigma}_{\bm{\bm{zz}}}\cdot\beta+\eta_{x}^{2}\\ \bm{\Sigma}(\Theta)_{xy}&:=&\Sigma(\Theta)_{\bm{z}x}\cdot\bm{\gamma}+\eta_{x}^{2}\times\theta+\eta_{xy}\\ \bm{\Sigma}(\Theta)_{\bm{z}y}&:=&\bm{\Sigma}_{\bm{zz}}\cdot\bm{\gamma}+\bm{\Sigma}(\Theta)_{\bm{z}x}\times\theta\\ \bm{\Sigma}(\Theta)_{yy}&:=&\bm{\gamma}\cdot\bm{\Sigma}_{\bm{zz}}\cdot\bm{\gamma}+2\times\bm{\gamma}\cdot\Sigma_{zx}(\Theta)\times\theta+\\ &&\theta^{2}\times\Sigma(\Theta)_{xx}+2\times\theta\times n_{xy}+\eta_{y}^{2}.\\ \end{array}$
@@ B.3.1 Usage
Like MASSIVE [bucur2020], this is a full Bayesian approach that returns a full posterior on $\theta$. Unlike MASSIVE, which is designed for soft sparsity constraints, our prior on $\gamma$ is on a Gaussian distributed disc with the radius being given a uniform prior on $[0,\tau]$ so that it is the closer match to the principle behind our hard constrained optimization method.

### v2

@@ 5 Related Work
Violations of the exclusion restriction are well-documented in genetics [Hemani et al., 2018] and econometrics [Berkowitz et al., 2008]. One strategy for estimating causal effects in such settings is to permit a large number of potentially invalid instruments under the assumption that their average bias will tend to zero in the limit [Bowden et al., 2015, Kolesár et al., 2015]. With weak monotonicity constraints, these approaches can also provide nonparametric bounds on local average treatment effects [Flores and Flores-Lagunes, 2013].
@@
Another family of methods starts from the assumption that some proportion of candidate instruments are valid and uses statistical procedures to focus on just those variables that satisfy (A1)-(A3). This can be achieved, for instance, via goodness of fit tests [Chu et al., 2001]; $L_{1}$-penalized regression for feature selection [Kang et al., 2016, Guo et al., 2018, Windmeijer et al., 2019]; independence tests for collider bias [Kang et al., 2020]; or modal validity assumptions in linear [Hartwig et al., 2017] and nonlinear [Hartford et al., 2021] IV models. Alternatively, data from multiple instruments can be pooled into a single variable using dimensionality reduction techniques [Kuang et al., 2020].
@@
There is a substantial literature on Bayesian approaches to causal inference in IV settings. Lenkoski et al. [2014] use Bayesian model averaging to select IVs based on the strength of their association with the treatment, as codified by the relevance criterion (A1). Shapland et al. [2019] extend this method to account for linkage disequilibrium in Mendelian randomization experiments, which violate the no confounding condition (A2). More recently, several authors have proposed spike-and-slab priors to select genetic variants in the face of horizontal pleiotropy [Bucur et al., 2020, Gkatzionis et al., 2021], thereby addressing (A3).
@@
Conley et al. [2012] propose ATE bounding methods given various kinds of prior information on leaky coefficients $\bm{\gamma}$, including a range restriction and a (non-uniform) prior distribution. For the former, they return a union of confidence intervals resulting from a grid of possible values for $\bm{\gamma}$. This method scales exponentially with $d_{\bm{Z}}$ and is potentially conservative as grid resolution grows finer. By contrast, our convex optimization approach is an efficient one-shot procedure that provides provably sharp bounds—in closed form, for the $L_{2}$ case. Their Bayesian proposal is similar to the method we compare against in Fig. 6.
@@
The literature on tetrad constraints in linear SEMs goes back over a century [Spearman, 1904], although the method was revived and refined following the publication of Spirtes et al. [2000]’s tetrad representation theorem. Our exclusion test builds on generalized results developed by numerous authors [Shafer et al., 1995, Sullivant et al., 2010, Spirtes, 2013], although we are to our knowledge the first to propose a Monte Carlo inference procedure for testing such claims.
@@
Partial identification intervals for counterfactual quantities can be computed exactly for discrete variables by formulating the problem as a polynomial program [Zhang et al., 2022, Duarte et al., 2023]. Though continuous data can always in principle be discretized with arbitrary precision, this quickly becomes intractable in the response function framework, as it leads to an exponential explosion of parameters. Some continuous alternatives have been proposed with applications to IV models [Kilbertus et al., 2020, Hu et al., 2021, Padh et al., 2023]. However, the neural architectures underlying these models can be notoriously unstable, and are ill-suited to the linear SEM setting, which is standard in much biological and econometric research.
@@ 6 Discussion
We have assumed in this paper that treatment effects are homogeneous throughout the population. However, a great deal of recent literature in causal machine learning has focused on heterogeneous treatment effects, where potential outcomes are presumed to vary as a function of pre-treatment covariates [Chernozhukov et al., 2018, Künzel et al., 2019, Nie and Wager, 2021]. Some authors have brought this framework into IV models, showing that tighter ATE bounds are possible with the help of observed confounders [Cai et al., 2007, Hartford et al., 2021, Levis et al., 2023]. Future work will consider conditional bounding methods, where extrema for $\theta$ may depend on instruments and/or other features causally antecedent to $X$.
A final note is that our method relies on user specification of the hyperparameter $\tau$. Ill-chosen thresholds may lead to issues, either in the form of overly conservative bounds (if $\tau$ is too high) or no bounds at all (if $\tau$ is below the minimum consistent with the data). In the worst case, invalid bounds will result from selecting a threshold that is high enough to satisfy the partial identification criterion but underestimates the true value of $\lVert\bm{\gamma}\rVert_{p}$. However, we emphasize that $\tau$-exclusion is a strictly weaker assumption than the classical exclusion criterion (A3), which is widely applied—rightly or wrongly—in IV analyses. In many cases, setting $\tau=0$—i.e., assuming perfectly exclusive instrumental variables—is an essentially arbitrary choice, and a tempting one given the veneer of certainty provided by obtaining a point estimate for the ATE. If nothing else, we hope to make practitioners think twice before falling back on this familiar default. Background knowledge is regularly used to guide hyperparameter selection, and it is reasonable to assume that in many applications practitioners will have an a priori sense of how much information leakage is likely between $\bm{Z}$ and $Y$. In particular, we can interpret this as a type of sensitivity analysis that asks how large $\tau$ can be such that the bounds exclude zero or effects of particular magnitude, and inquire with a practitioner whether larger values of $\tau$ are scientifically plausible.
@@ 7 Conclusion
Acknowledgments
RS has been funded by the EPSRC Fellowship EP/W024330/1 and the EPSRC AI Hub EP/Y00759X/1.
@@ A.1 Proof of Lemmas 1 and 2
Consider Eqs. 1, 2, and 3, which define our model. Evaluating the product of $X$ and $\bm{Z}$ gives a relationship between covariances:
@@ A.3 Proof of Lemma 4
where $\theta^{*},\bm{\gamma}^{*}$ denote the true unobservable parameters. By the convexity of the norm, the solutions to Eq. 12 for any $\tau\in[\check{\tau}_{p},\tau^{*}_{p})$ will fail to capture the true ATE, as the resulting horizontal line lies below the point $(\lVert\bm{\gamma}^{*}\rVert_{p},\theta^{*})$. Even a $\bm{\Sigma}$-oracle—who, recall, has access to the population covariance matrix but not the latent parameters $\theta^{*},\bm{\gamma}^{*}$—will return invalid bounds if queried with a threshold in this half-closed interval. For this reason, we call this band the error region.
@@
We have thus far assumed that the oracle threshold strictly exceeds the theoretical minimum, but these parameters may coincide. If $\tau^{*}_{p}=\check{\tau}_{p}$, then the error region is empty and all identifiable bounds are valid. Moreover, if $g_{p}$ attains a unique minimum—as it must for all strictly convex norms, i.e. $p\in(1,\infty)$—then there exists just a single solution to Eq. 12 for $\tau=\tau^{*}_{p}=\check{\tau}_{p}$. In this case, lower and upper bounds for the ATE coincide and the causal parameter is point identified as $\theta^{*}=\check{\theta}_{p}$. Note that this fortuitous circumstance occurs with Lebesgue measure zero, as it imposes a nontrivial polynomial constraint on covariance parameters [Okamoto, 1973].
@@
For every $P_{\bm{\Sigma}}\in\mathcal{M}$, there exists some nearest null neighbor $Q^{*}_{\bm{\Sigma}}\in\mathcal{M}_{0}$ (not necessarily unique) satisfying
@@
for all $u\in[0,1]$ [Lehmann and Romano, 2005]. In other words, $p_{\text{MC}}$ is uniformly distributed under $H_{0}$, as desired, and the Monte Carlo null distribution has converged on the target $G_{n}^{\bm{\Sigma}^{0}}$. We add one to the numerator and denominator as a necessary finite sample adjustment.
@@
It may not be immediately clear that bootstrapping is appropriate for ATE bounds. After all, it is well known that the bootstrap cannot provide a valid sampling distribution for fixed order statistics such as ranks, or a target parameter that lies on the boundary of the parameter space [Andrews, 2000]. But though we refer to our bounds as “min” and “max” solutions, that does not mean they are calculated via fixed order statistics. On the contrary, each represents a continuous solution to a differentiable optimization task, not the smallest or largest element in a discrete set. The only boundary condition for our target parameters is $\theta^{-}\leq\theta^{+}$, which automatically holds under the partial identifiability criterion $\tau\geq\check{\tau}_{p}$. As our estimator is undefined for samples that violate the criterion, our sampling distribution is always conditioned on this event.
In general, any statistic that is a differentiable function of sample moments admits an Edgeworth expansion and can therefore have its distribution consistently estimated via bootstrap resampling [Hall, 1992, Davison and Hinkley, 1997]. Recall that our ATE bounds represent the intersection of (a) a $\tau$-feasible region that is fixed a priori; and (b) the $L_{p}$ norm of $\bm{\gamma}=\bm{\alpha}-\theta\bm{\beta}$, where the latter two vectors are defined as dot products of covariance parameters. Resulting bounds vary smoothly under resampling, since $\bm{\alpha}$ and $\bm{\beta}$ are differentiable with respect to $\bm{\Sigma}$. Though more generalized relaxations of the exclusion criterion may introduce discontinuities or other issues, our formulation of $\tau$-exclusion poses no such difficulties. The resulting bootstrap distributions are asymptotically valid and practically useful, providing statistical inference without any parametric assumptions.
Of course, it is perfectly possible that some bootstrap samples may have to be discarded if the intersection of regions (a) and (b) is empty. In such cases, we simply restrict attention to those bootstraps that satisfy the partial identifiability criterion $\tau>\check{\tau}_{p}$, which should represent a non-negligible proportion of all bootstraps if the inequality is satisfied in the original dataset. This procedure is akin to sampling under a feasibility condition, and requires no extra steps to maintain bootstrap consistency, as in Andrews [2000] or Ramsahai and Lauritzen [2011].
@@ B.3 Bayesian baseline
$\begin{array}[]{rcl}\bm{\Sigma}(\Theta)_{\bm{zz}}&:=&\ \bm{\Sigma}_{\bm{zz}}\\ \bm{\Sigma}(\Theta)_{\bm{z}x}&:=&\ \bm{\Sigma}_{\bm{zz}}\beta\\ \bm{\Sigma}(\Theta)_{xx}&:=&\bm{\beta}\cdot\bm{\Sigma}_{\bm{zz}}\cdot\beta+\eta_{x}^{2}\\ \bm{\Sigma}(\Theta)_{xy}&:=&\Sigma(\Theta)_{\bm{z}x}\cdot\bm{\gamma}+\eta_{x}^{2}\times\theta+\eta_{xy}\\ \bm{\Sigma}(\Theta)_{\bm{z}y}&:=&\bm{\Sigma}_{\bm{zz}}\cdot\bm{\gamma}+\bm{\Sigma}(\Theta)_{\bm{z}x}\times\theta\\ \bm{\Sigma}(\Theta)_{yy}&:=&\bm{\gamma}\cdot\bm{\Sigma}_{\bm{zz}}\cdot\bm{\gamma}+2\times\bm{\gamma}\cdot\Sigma_{zx}(\Theta)\times\theta+\\ &&\theta^{2}\times\Sigma(\Theta)_{xx}+2\times\theta\times n_{xy}+\eta_{y}^{2}.\\ \end{array}$
@@ B.3.1 Usage
Like MASSIVE [Bucur et al., 2020], this is a full Bayesian approach that returns a full posterior on $\theta$. Unlike MASSIVE, which is designed for soft sparsity constraints, our prior on $\gamma$ is on a Gaussian distributed disc with the radius being given a uniform prior on $[0,\tau]$ so that it is the closer match to the principle behind our hard constrained optimization method.